# Visual Scratchpads: Enabling Global Reasoning in Vision

### v1

@@ 2.1 Globality, locality and scratchpads in the visual domain
Figure 5: Comparison between training from scratch and initializing with a pre-trained model, while varying the task complexity. Pre-training is not sufficient to guarantee convergence for complex tasks. Convergence without pre-training is not possible even for easier tasks.
@@
Next, we discuss the connection of visual scratchpads to scratchpads used in text (Nye et al., 2021) and why it helps. The idea of scratchpad generally refers to training the models with intermediate reasoning steps, so that they generate both the reasoning steps and the final answer during inference. For instance, consider math questions with simple numerical answers. Nye et al. (2021) showed that training language models to first output the intermediate steps of the solution and then the final numerical answer results in superior accuracy than training the model to directly output the answer.
Abbe et al. (2024) have shown that scratchpads can reduce the globality degree and by doing so reduce the learning complexity. More specifically, the scratchpad can provide intermediate targets $Y_{1},\ldots,Y_{m}$ such that $Y_{m}=Y$ and each $Y_{i}$ is of low globality given the previous intermediate targets and the input which makes predicting them in a sequential manner easily learnable. The same is true for our proposed datasets and multiple-frame scratchpads as each scratchpad frame is a low globality function of the previous frame and also the label is a low globality function of the final frame. For example, in each frame of the multiple-frame scratchpad for the cycles task at most two nodes are colored, and determining these nodes is possible by using a few patches. Similarly, determining the final label from the final scratchpad frame is a local operation since checking one node’s color provides significant information on the label. Note that to compute the label, one has to check whether all nodes are colored or not. However, even checking whether one node is colored or not provides non-trivial information on the output even though it may not determine the output perfectly (if that node is colored). Thus, the multi-frame scratchpad can reduce the learning complexity of tasks by breaking their globality degree. Interestingly, the single-frame scratchpad model can also break the globality degree of the task. As explained before, predicting the label from the complete scratchpad is a local function. Note that learning the full scratchpad from the input image is also a low globality function. To see this, note that coloring the first three nodes of the scratchpad (which corresponds to non-trivial mutual information) is a local function as it requires a limited number of patches which allows weak learning of the scratchpad. A particularly interesting phenomenon is that this weak learning can result in strong learning of the scratchpad frame through hierarchical learning. We explain this in more details in Section 4.4.
3 Methodology
As discussed in the introduction, we have three supervision formats for our visual reasoning datasets: no scratchpad, single-frame scratchpad, and multi-frame scratchpad. We use a vision Transformer (ViT) (Dosovitskiy et al., 2020) backbone for all supervision modes.
@@ 3 Methodology
No scratchpad baseline. In this case, the image is given to the model, and the model has to produce the label directly. We use a ViT architecture with a classification token, CLS, for this setting. We use a linear layer on the CLS token features to compute the label logits. We use the cross-entropy loss function on the logits for training the model. As it will be shown in Section 4.1, this model is not capable of learning the proposed datasets. Hence, we next introduce the single-frame scratchpad.
3.1 Single-frame scratchpad
@@
For generating the scratchpad in a single-frame format, we make some modifications to the no-scratchpad architecture above. Here, the model predicts both the label and the scratchpad (the complete one in the single-frame format). We keep the ViT encoder with a CLS as the backbone and add a linear layer to the hidden representation of the last transformer layer to predict the scratchpad image. During training, we use cross-entropy loss to supervise the label and pixel-wise mean-squared loss similar to He et al. (2022); El-Nouby et al. (2024) to supervise the scratchpad image. In Section 4.1, we show that the single-frame model has better performance than the no-scratchpad model, and for large enough models it may be able to learn the proposed tasks. In Section 4.2, we evaluate the out-of-distribution (OOD) performance of this model. In the next part, we introduce our inductive model for multi-frame scratchpads which has superior performance in both the in-distribution and OOD settings.
3.2 Multi-frame scratchpad and Inductive scratchpad model
@@
In this section, we introduce the inductive scratchpad model that is used for the multi-frame scratchpad setting where the model predicts all the scratchpad frames in a sequential and Markovian manner. In this setting, the model predicts each scratchpad frame based only on the previous scratchpad frame predicted by the model (or the input at the beginning). For example, the model first predicts the first scratchpad frame using the input image. Then, it uses the first frame to predict the second one, and so on. More precisely, the model has a recurrent component $\mathcal{M}$ that takes an input image (either the input image or a scratchpad frame) as input and predicts three outputs: the next scratchpad frame ($\hat{f}$), the label ($\hat{y}$), and a binary variable for halting ($\hat{h}$). This recurrent module is applied to the input image and the subsequent intermediate frames until the halting signal activates (or an upper limit of recurrences is reached). The predicted label at the last recurrence is the predicted label of the model. Note that generating each scratchpad frame depends only on the last generated frame (or the input image) and there is no part of the recurrent module to record the history. As a result, the model is independent of the number of scratchpad frames used in each sample.
@@
Training procedure. During training, we initially used teacher forcing, which consists of providing the model with perfect frames from the training set. However, this approach creates a discrepancy during inference, where the model sees its own generated frames as input. Generated frames may exhibit a slightly different distribution as the reconstructions are not perfectly accurate. While this issue is well studied in text generation, where discrete tokens are used, it becomes more pronounced in vision tasks with continuous outputs. To mitigate this discrepancy, we use an alternated training procedure where the model sees perfect frames 50% of the times, and generated frames the other 50%. This ensures that the model learns to handle imperfect inputs, leading to improved performance during inference. More details on this training procedure can be found in Appendix B.1.
4 Experiments
@@
In this section, we show the performance of different methods on our proposed datasets focusing on required model size and OOD generalization. Each of our datasets contains $10^{6}$ (1M) training samples. See Appendix A for more details on the experiments.
@@
Figure 6: Validation accuracy for different datasets learned by different methods and model sizes. We can see that the model without a scratchpad is not capable of learning any of these tasks, while for large enough models, the single-frame scratchpad model may be able to learn. Further, the inductive scratchpad model can learn all the tasks with smaller models than the single-frame scratchpad model.
4.1 Model size experiments
First, we compare the performance of different methods with varying model sizes on our proposed datasets. In particular, we compare the no scratchpad baseline, the single-frame scratchpad model, and the inductive scratchpad model used for multi-frame scratchpad prediction. Moreover, we use four different sizes for the ViT encoder of our models: small, base, large, and huge, which have respectively around $22M$, $86M$, $307M$, and $632M$ parameters (see Appendix B for detailed specifications). The accuracy of different methods with different model sizes is shown in Figure 6. It can be seen that the no-scratchpad baseline is not able to go beyond random accuracy for any of these tasks. On the other hand, the single-frame scratchpad model can learn the cycles 24 and maze (rect.) 32 tasks for large enough models while it still cannot learn the strings 20 task. The inductive scratchpad model used for the multi-frame prediction, however, learns all the proposed tasks even with smaller models. We report the results for the circular maze dataset in Appendix D.
We note that the number of parameters for the three methods is very similar. The single-frame model only adds a linear layer for predicting the scratchpad to the no-scratchpad baseline. Likewise, the inductive scratchpad model only adds a linear layer for predicting the halt signal to the single-frame model. However, during inference, the inductive scratchpad model is applied a variable number of times. Hence, the inductive scratchpad model uses compute adaptively depending on the complexity of the sample, and therefore its inference time compute is usually larger than the baselines.
@@ 4.2 OOD generalization
Next, we consider the out-of-distribution (OOD) generalization performance of different methods where we show the inductive scratchpad model has a superior OOD generalization performance. This observation is due to the fact that the inductive scratchpad model only learns the steps of the reasoning process, and as a result, is independent of the number of reasoning steps required allowing it to generalize to harder problems with its adaptive compute time.
Figure 7: OOD experiments where the model is trained on cycles 12 and tested on more complex cycles tasks.
Table 1: OOD performance for maze 24 (rectangular) dataset. While both models are good in-distribution, the inductive scratchpad achieves almost perfect accuracy on OOD while the single-frame model hardly goes beyond the random baseline (50%).
Method Accuracy (%)
@@
Single-frame 100.0 54.4
Inductive 99.8 99.8
For the cycles task, we can control the complexity of the task with the number of nodes. In particular, for OOD experiments, we consider training on samples with 12 nodes and then testing on samples with a higher number of nodes and thus higher complexity. The results are visualized in Figure 7.
For the maze tasks, we do not change the size of the maze as it would result in resolution inconsistencies. Instead, we create a dataset of easier samples for training (e.g., if the source and sink points are connected their distance is less than or equal to $30$) and use the main task dataset for validation. We explain the OOD training datasets for the maze tasks in more detail in Appendix C. The OOD results for the rectangular maze task are shown in Table 7. It can be seen that the inductive scratchpad model achieves almost perfect accuracy on OOD samples while the single-scratchpad model performs slightly better than random. We present more OOD experiments in Appendix D.
@@ 4.3 Ablations
Method ID (%) OOD (%)
Inductive 99.99 88.21
Inductive (only TF) 99.91 85.15
Inductive w/o halting 99.94 80.89
Multi-frame 99.99 64.83
Single-frame 99.95 64.79
Table 2: Comparison of in-distribution and average out-of-distribution accuracy for multi-frame and single-frame scratchpad variants.
The success of the inductive scratchpad model can be attributed to several factors such as increased supervision during training, halting mechanism, and the combination of teacher forcing (TF) and training on the output distribution of the model at training. In this part, we provide experiments that show the effect of these elements. To do this, we consider a multi-frame baseline model which is similar to the single-frame model with the difference that it has multiple heads for predicting multiple scratchpad frames. [^5] We use a preset number of heads and repeat the last scratchpad frame for heads with a missing frame. For the “Inductive w/o halting” baseline we simply set a fixed large number of steps. The OOD performance of different variations on the cycles task is reported in Table 2. It can be seen that removing aspects such as teacher forcing and adaptive (early) halting can each cause minor reductions in the performance of the inductive scratchpad model, whereas, the multi-frame baseline provides no performance gain on OOD samples compared to the single-frame model. More ablations are shown in Appendix D.4.
@@ 4.4 Staircase learning phenomenon
In our experiments with generating a single-frame scratchpad, we observed progressive hierarchical learning over the scratchpad image prediction task. Consider the cycles task as a running example. In the training set, we always color the cycle that passes the rightmost node. In the scratchpad generations, we can observe that the model first learns to color the rightmost node. Then it learns to color the two neighbors that are connected to the initial node. Similarly, at each of the later stages, it learns to color roughly two more nodes (from the two sides). See Figure 8 for a visualization.
@@
Figure 8: Generated scratchpads for an example at different stages during training. We have increased the contrast of the images for better visualization. It can be seen that the model first learns to color the rightmost node and then it goes one distance further each time during training.
@@
These hierarchical learning phenomena have been previously observed and proven in theoretical settings, in particular, in the context of learning sparse Boolean functions where it is known as the staircase behavior (Abbe et al., 2022, 2023a). The staircase phenomenon states that if the target function has some hierarchical structure and is composed of different parts with different difficulties, learning easier parts first can boost learning for the harder parts. To be more precise, assume we have $n$ i.i.d. uniform Boolean variables $x_{1},\ldots,x_{n}\in\{\pm 1\}$. It is well known that the difficulty of learning degree $k\leq n/2$ parity function, e.g., $x_{1}x_{2}\cdots x_{k}$ increases as $k$ increases. In particular, if $k=\omega_{n}(1)$ then learning the parity function is not possible in polynomial time with regular MLPs (and statistical query methods) and learning complexity of degree $k$ parity for constant $k$ scales with $n^{k}$ (Abbe & Sandon, 2023). However, Abbe et al. (2022) show that functions such as $x_{1}+x_{1}x_{2}+\cdots+x_{1}x_{2}\cdots x_{k}$ can be learned in $O(n)$ time. This is because the network can first learn the ‘easy’ linear component $x_{1}$. Now, for learning $x_{1}x_{2}$ the model no longer needs to find two variables (which would scale with $\binom{n}{2}$), but it needs to only find $x_{2}$ since it knows that $x_{1}$ is in the support and can navigate the search space more efficiently to learn the terms like $x_{1}x_{i}$.
@@
Considering the cycles task in the single-frame scratchpad again, coloring the first three nodes is a low globality function and can be learned easily. Next coloring the next two nodes once the coloring of the first three nodes is learned is a low globality function (similar to $x_{1}\cdots x_{i}$ when $x_{1}+x_{1}x_{2}+\cdots x_{1}\cdots x_{i-1}$ is learned). More precisely, define $Y_{k}$ to be the color of all nodes (and edges) with a distance less than or equal to $k$ from the rightmost node ($12k+$ 1 nodes in total). $Y_{1}$ is a local target, moreover, coloring $Y_{k+1}$ correctly once $Y_{k}$ is learned is of constant globality degree. This staircase structure allows the model to learn $Y_{1},Y_{2},\ldots$ and finally the complete scratchpad frame sequentially during training as observed in Figure 8. Note that in the example of the cycles task the intrinsic staircase structure of the single-frame scratchpad coincided with the multi-frame scratchpad, however, that is not necessarily always the case.
@@
This example shows that the globality degree does not satisfy the triangle inequality. In other words, we show that for input $X$ and target $Y$ and diverging globality degree (e.g., increasing number of nodes in the cycles task), there exists a single-frame scratchpad $X_{1}$ such that globality degrees of $X_{1}$ from $X$ and $Y$ from $X_{1},X$ is constant. Thus, a single-frame scratchpad can make both efficient weak and strong learning (through the staircase effect) possible.

### v2

@@ 2.1 Globality in the visual domain
Figure 5: Comparison between training from scratch and initializing with a pre-trained model, as the task complexity varies. Pre-training is not sufficient to guarantee convergence as the complexity increases. Learning without pre-training is not possible even for simple instances.
We stress that a model trained from scratch was not able to learn the cycles task even when no patch was masked. To better explain this phenomenon, in Figure 5, we compare the performance when initializing with a pre-trained model (Fang et al., 2023) versus a model trained from scratch on the cycles task of varying size. This shows that as the task complexity increases with the number of nodes, none of the models is able to learn the cycles task, meaning that even strong internet-sourced priors in the pre-trained model are not helpful.
2.2 Global visual tasks are hard for multi-modal LLMs
In this section, we assess the difficulty of our proposed tasks for state-of-the-art multi-modal large language models (LLMs). Specifically, we selected two of our tasks for this evaluation: the cycles task and the rectangular maze [^5] For the rectangular maze task, we slightly reduced the size of the blue and red cells in our images so that they do not touch the walls of the maze, eliminating any potential ambiguities for the model.. These tasks are relatively natural and likely resemble examples encountered during pretraining. Moreover, they are easier than their respective counterparts, i.e., the strings task and the circular maze. [^6] The models can potentially come up with symbolic representations for the cycles and rectangular maze tasks, whereas doing so for the strings and circular maze tasks is more challenging.
For evaluation, we used the GPT-4o (Hurst et al., 2024), o3, and o4-mini (OpenAI, 2025) models and 1000 test samples per task (500 from each class). We considered two settings: (1) providing the model with a detailed prompt and asking it to solve the task, and (2) using a similar prompt along with four in-context examples (two from each class) presented as images. We disabled tool calling in both settings as our main interest is to measure the model’s visual reasoning capabilities without allowing it to execute any code. The results are shown in Table 1. We observe that models perform barely better than random as the tasks get harder. Interestingly, in-context learning (ICL) does not appear to be particularly beneficial in our setting. See Appendix C.5 for more details and the prompts used.
Task Size GPT-4o o4-mini o3
Prompt ICL Prompt ICL Prompt ICL
Cycles 6 nodes $61.9\%$ $58.9\%$ $68.9\%$ $62.2\%$ $63.4\%$ $62.1\%$
8 nodes $51.6\%$ $50.4\%$ $55.4\%$ $52.9\%$ $54.7\%$ $51.4\%$
10 nodes $49.9\%$ $48.8\%$ $50.5\%$ $48.4\%$ $49.8\%$ $51.3\%$
Maze (rect.) $4\times 4$ $50.6\%$ $55.4\%$ $55.3\%$ $52.4\%$ $52.4\%$ $51.9\%$
$6\times 6$ $48.4\%$ $51.8\%$ $49.4\%$ $51.3\%$ $51.2\%$ $48.6\%$
Table 1: The accuracy of different models on our proposed tasks. Prompt columns represent the experiments where no in-context example was included while ICL is used for experiments with 4 in-context examples. Note that $50\%$ would correspond to the accuracy of a random model. Also, we have used 1000 samples for each task, giving us a 95% confidence interval of approximately ±$\pm 3\%$.
2.3 Breaking globality with chain-of-sketch
@@
We next discuss the connection between our chain-of-sketch and the scratchpad and chain-of-thought (CoT) used in text and why it helps. The idea of CoT generally refers to training the models with intermediate steps, so that they generate both the reasoning steps and the final answer at inference. For instance, Nye et al. (2021) showed that for simple math questions training language models to output intermediate steps before the final answer leads to higher accuracy than training the model to directly output the answer. Wei et al. (2023) further developed this methodology and showed that pre-trained models can generate these intermediate steps with in-context examples and prompting as well. See related work in Section 5 for further discussion.
In (Abbe et al., 2024), it is shown that chain-of-thought can reduce the globality degree and, in doing so, the learning complexity as well. More specifically, chain-of-thought can provide intermediate targets $Y_{1},\ldots,Y_{m}$ such that $Y_{m}=Y$ is the final target and each $Y_{i}$ is of low globality given the previous intermediate targets and the input, which makes predicting them in a sequential manner easily learnable. The same is true for our proposed datasets and CoS as each sketch frame is a low globality function of the previous frame and the label is also given by a low globality function of the final frame. For instance, each frame of our CoS in the cycles task reduces the globality degree from a quantity growing with the number of nodes to a constant, as in each step only the neighboring nodes are considered.
Single-frame CoS and hierarchical learning. The single-frame CoS collapses all the steps of the multi-frame CoS (into the final frame). Both CoS, however, break the globality of the original target. This may not be obvious for the single-frame CoS, but it results from the fact that there are parts of the single-frame CoS target that can be predicted with low globality from the input image, i.e., weak learning of this single sketch is possible. For instance, for the cycles task, one always starts by coloring the edges adjacent to the preset vertex. The model can then learn how to progress the coloring ‘on its own’, rather than relying on the guidance of a multi-frame CoS, and thereby turns its weak learning into strong learning. This hierarchical learning mechanism is also known as the staircase phenomenon (Abbe et al., 2022b, 2023a), and is further discussed in Section 4.4. Note that the label is also a low globality function of the full sketch in the single-frame CoS.
3 The CoS methodologies
A CoS consists of adding intermediate visual steps to learn a given task. The goal is generally to have a sequence having a reduced globality compared to the original target’s globality. In this paper, we consider three variants of CoS to evolve from the no-CoS model to break the globality of the target: single-frame CoS, multi-frame CoS, and inductive CoS. We describe these below and discuss similarities with previous work in Sections 5.2, 5.3, and 5.4.
No-CoS baseline. This baseline corresponds to predicting the target label directly from the input image without any intermediate step.
@@ 3 The CoS methodologies
We use a ViT (Dosovitskiy et al., 2020) architecture with a classification token, CLS, for this setting. We use a linear layer on the CLS token features to compute the label logits and we use the cross-entropy loss function for training. As shown in Section 4.1, this model is not capable of learning the proposed datasets.
3.1 Single-frame CoS
In this case, the CoS introduces a single frame to be predicted as the intermediate target. For instance, for the cycles task it would be coloring the cycle that passes through the rightmost node, while for the maze task, it would be coloring the region reachable from the start cell of the maze. See Figure 2 for examples of the single-frame CoS method for different tasks.
@@ 3.1 Single-frame CoS
For implementation, we keep the ViT encoder with a CLS as the backbone and add a linear layer to the hidden representation of the last Transformer layer to predict the sketch image. During training, we use cross-entropy loss to supervise the label and a pixel-wise mean-squared loss similar to He et al. (2022); El-Nouby et al. (2024); Fini et al. (2024) to supervise with the sketch image. In Section 4.1, we show that the single-frame CoS improves accuracy over the no-CoS model, and for large enough models, it may be able to learn the proposed tasks.
3.2 Inductive CoS
In CoS with multiple frames, the intermediate target of the single-frame CoS is decomposed into multiple steps. Namely, rather than providing the full CoS image in one single frame, the model is provided with a sequence of frames that gradually lead to the full CoS image. For example, for the cycles task each CoS frame corresponds to coloring two edges at a time (incrementing the distance to the preset vertex). See Appendix C for the details of CoS frames and Appendix F.1 for examples of CoS frames for different tasks.
One way to implement CoS with multiple frames is to use a similar architecture to the single-frame CoS and predict all the intermediate images in parallel. We use “multi-frame CoS” to refer to this preliminary model. This model serves as a useful baseline for ablation studies in Section 4.3. Next, we introduce the inductive CoS method.
@@ 3.2 Inductive CoS
Inductive CoS has also multiple steps/frames, but these are learned sequentially with an inductive model. Namely, the sequence is Markovian, i.e., each new frame depends only on the previous one. More precisely, the model has a recurrent component $\mathcal{M}$ that takes an input image (either the input image or a CoS frame) and predicts three outputs: the next CoS frame ($\hat{f}$), the label ($\hat{y}$), and a binary halting variable ($\hat{h}$). This recurrent module is applied to the input image and the subsequent intermediate frames until the halting signal is triggered (or an upper limit of recurrences is reached). The predicted label at the last recurrence is the predicted label of the model. Note that generating each CoS frame depends only on the last generated frame (or the input image) and the recurrent module does not retain a history of prior frames. As a result, the model is independent of the number of sketches (frames) used in each sample.
We will see that the inductive CoS affords better OOD generalization than the single-frame and multi-frame CoS, due to its compositional structure that is independent of the number of frames a CoS has. The inductive CoS will also enable smaller models to learn when the other CoS variants do not on most considered tasks.
@@
Training procedure. For training, we initially used teacher forcing, which consists of providing the model with perfect frames from the training set. However, this approach creates a discrepancy during inference, where the model sees its own generated frames as input. Generated frames may exhibit a slightly different distribution as the reconstructions are not perfectly accurate. While this issue is well studied in text generation, where discrete tokens are used, it becomes more pronounced in vision tasks with continuous outputs. To mitigate this discrepancy, we use an alternated training procedure where the model sees perfect frames 50% of the time, and generated frames the other 50%. This ensures that the model learns to handle imperfect inputs, leading to improved performance during inference.
4 Experiments on OOD generalization and model size
@@
In this section, we show the performance of different methods on our proposed datasets, focusing on required model size and OOD generalization. Each of our datasets contains $10^{6}$ (1M) training samples. See Appendix A for more details on the experiments. Further see Table 1 in Section 2.2 for results with multi-modal LLMs such as GPT-4o (Hurst et al., 2024), o3, and o4-mini (OpenAI, 2025) where we show these models perform only slightly better than random on our proposed tasks.
4.1 Model size experiments
First, we compare the performance of different methods with varying model sizes on our proposed datasets. In particular, we compare the no-CoS baseline, the single-frame CoS, and the inductive CoS model used for multi-frame sketch prediction. Moreover, we use four different sizes for the ViT (Dosovitskiy et al., 2020) encoder of our models: small, base, large, and huge, which have respectively around $22M$, $86M$, $307M$, and $632M$ parameters (see Appendix B for detailed specifications). The accuracy of different methods with different model sizes is shown in Figure 6. It can be seen that the no-CoS baseline is not able to go beyond random accuracy for any of these tasks. On the other hand, the single-frame CoS model can learn the PVR($7\times 7,k=4$), cycles 24, maze (circ.) 16, and maze (rect.) 32 tasks for appropriately large models while it still cannot learn the strings 20 task. The inductive CoS model used for the multi-frame prediction, however, learns the proposed tasks even with smaller models. We report the results for the circular maze and PVR dataset in Appendix D.
Compute overhead. We note that the number of parameters for the three methods is very similar. The single-frame model only adds a linear layer for predicting the CoS to the no-CoS baseline. Likewise, the inductive CoS model only adds a linear layer for predicting the halt signal to the single-frame model. However, during inference, the inductive CoS model is applied a varying number of times and uses compute adaptively depending on the complexity of the sample, and therefore usually requires a higher compute during inference compared to the single-frame CoS for a model of the same size. Nevertheless, as shown in Figure 6, we can often use much smaller models with the inductive CoS model. We further discuss compute trade-offs in Appendix B.2.
@@ 4.1 Model size experiments
Figure 6: Validation accuracy for different datasets learned by various methods and model sizes. We can see that the model without a CoS is not capable of learning any of these tasks, while for large enough models, the single-frame CoS model may be able to learn. Further, the inductive CoS model can learn all the tasks with smaller models than the single-frame CoS model.
@@ 4.2 OOD generalization
Next, we consider the out-of-distribution (OOD) generalization performance of different methods and show that the inductive CoS model exhibits superior OOD generalization. This observation is due to the fact that the inductive CoS model only learns the steps of the reasoning process, and as a result, is independent of the number of reasoning steps required—allowing it to generalize to harder problems with its adaptive compute time.
Figure 7: The model is trained on Cycles 12 and tested on more complex instances.
Dataset Method Accuracy (%)
@@
Maze 24 (Rect.) single-frame CoS 100.0 54.4
inductive CoS 99.8 99.8
PVR Grid 7x7 single-frame CoS 100.0 48.6
inductive CoS 99.5 82.2
Table 2: Out-of-distribution (OOD) performance comparison for maze 24 (rectangular) and PVR grid 7x7 datasets. The inductive CoS demonstrates superior OOD generalization compared to the single-frame CoS across both tasks.
For the cycles task, we control task complexity via the number of nodes. In particular, for OOD experiments, we consider training on samples with 12 nodes and then testing on samples with a higher number of nodes and thus higher complexity. The results are visualized in Figure 7. For the maze tasks, we keep the maze size the same. Instead, we create a dataset of easier samples for training (e.g., if the source and sink points are connected their distance is less than or equal to $30$) and use the main task dataset for validation. We explain the OOD training datasets for the maze tasks in more detail in Appendix C. The OOD results for the rectangular maze task are shown in Table 2. It can be seen that the inductive CoS model achieves almost perfect accuracy on OOD samples while the single-frame model performs slightly better than random. For the PVR task, we use a $7\times 7$ grid and $k=2,3,4$ images per row for training and $7\times 7$ grid and $k=5$ images per row for OOD samples. As shown in Table 2, the inductive model achieves $\approx 80\%$ accuracy while the OOD performance of the single-frame model is at chance level. We present more OOD experiments in Appendix D.
Method ID (%) OOD (%)
inductive CoS 100.0 88.2
inductive CoS (only TF) 99.9 85.2
inductive CoS w/o halting 99.9 80.9
multi-frame CoS 100.0 64.8
single-frame CoS 100.0 64.8
Table 3: Comparison of in-distribution and average OOD accuracy for CoS variants on the cycles task.
@@ 4.3 Ablations
The success of the inductive CoS model can be attributed to several factors, such as increased supervision during training, the halting mechanism, and the combination of teacher forcing (TF) and training on the model’s own output distribution. In this section, we use the cycles task as a reference and provide extensive ablation experiments to quantify the contribution of each component. The computational implications of these components are also discussed in Appendix B.2.
Supervision. To show the importance of how supervision is applied—as well as its extent—we compare the inductive and multi-frame CoS models. While both methods have the same amount of target supervision, the inductive CoS exhibits much stronger OOD generalization, as shown in Table 3 (see also Appendix D.5).
Teacher forcing. Here, we assess how much our improved training procedure contributes to performance. Specifically, we consider a baseline, “inductive CoS (only TF),” which uses the standard training procedure: it relies solely on teacher forcing and is not trained on the distribution of generated frames. Compared to the full inductive CoS model, this variant suffers a $3\%$ drop in performance.
Halting. We design an “inductive CoS w/o halting” baseline where we simply set a fixed large number of steps, eliminating the need for a halting signal. Compared to the full model, we observe an $\approx 7\%$ improvement in OOD generalization when dynamic halting is used. While the gap is significant, it is clearly smaller than the difference between inductive and non-inductive methods.
@@ 4.4 Staircase learning phenomenon
In our experiments with the single-frame model, we observed progressive hierarchical learning over the CoS image prediction task. Consider the cycles task as a running example. In the training set, we always color the cycle that passes the rightmost node. For sketch generations, we can observe that the model first learns to color the rightmost node. Then it learns to color the two neighbors that are connected to the initial node. Similarly, at each of the later stages of training, it learns to color roughly two more nodes (from the two sides). See Figure 8 for a visualization.
@@
Figure 8: Generated sketches for an example at different stages during training. We have increased the contrast of the images for better visualization. It can be seen that the model first learns to color the rightmost node and then it goes one distance further each time during training.
@@
These hierarchical learning phenomena have been previously observed and proven in theoretical settings, in particular, in the context of learning sparse Boolean functions where it is known as the staircase behavior (Abbe et al., 2022b, 2023a). The staircase phenomenon states that if the target function has some hierarchical structure and is composed of different parts with different difficulties, learning easier parts first can boost learning for the harder parts. To be more precise, assume we have $n$ i.i.d. uniform Boolean variables $x_{1},\ldots,x_{n}\in\{\pm 1\}$. It is well known that the difficulty of learning degree $k\leq n/2$ parity function, e.g., $x_{1}x_{2}\cdots x_{k}$ increases as $k$ increases. In particular, if $k=\omega_{n}(1)$ then learning the parity function is not possible in polynomial time with regular MLPs (and statistical query methods) and learning complexity of degree $k$ parity for constant $k$ scales with $n^{k}$ (Abbe and Sandon, 2023). However, Abbe et al. (2022b) show that functions such as $x_{1}+x_{1}x_{2}+\cdots+x_{1}x_{2}\cdots x_{k}$ can be learned in $\tilde{O}(n)$ time. This is because the network can first learn the ‘easy’ linear component $x_{1}$. Now, for learning $x_{1}x_{2}$ the model no longer needs to find two variables (which would scale with $\binom{n}{2}$), but it needs to only find $x_{2}$ since it has already learned that $x_{1}$ is in the support and can navigate the search space more efficiently to learn the terms like $x_{1}x_{i}$.
@@
Considering the cycles task in the single-frame CoS again, coloring the first three nodes is a low globality function and can be learned easily. Coloring the next two nodes once the coloring of the first three nodes is learned is a low globality function (similar to $x_{1}\cdots x_{i}$ when $x_{1}+x_{1}x_{2}+\cdots x_{1}\cdots x_{i-1}$ is learned). More precisely, define $Y_{k}$ to be the coloring of all nodes (and edges) with a distance less than or equal to $k$ from the rightmost node ($12k+$ 1 nodes in total). $Y_{1}$ is a local target, moreover, coloring $Y_{k+1}$ correctly once $Y_{k}$ is learned is of constant globality degree. This staircase structure allows the model to learn $Y_{1},Y_{2},\ldots$ and finally the complete sketch sequentially during training as observed in Figure 8.777For extra clarity, the visualization is enhanced in Figure 8. Nevertheless, this behavior is also evident in the non-enhanced outputs, as shown in the third row of Figure 22. Note that in the example of cycles task the intrinsic staircase structure of the single-frame model coincided with the CoS, however, that is not necessarily always the case.
@@
This example shows that the globality-degree does not satisfy the triangle inequality. In other words, we show that for input $X$ and target $Y$ and diverging globality degree (e.g., increasing number of nodes in the cycles task), there exists a single-frame sketch $X_{1}$ such that globality degrees of $X_{1}$ from $X$ and $Y$ from $X_{1},X$ is constant. Thus, a single-frame CoS can make both efficient weak and strong learning (through the staircase effect) possible.
Moreover, this hierarchical learning phenomenon is not limited to the cycles task. Figure 23 demonstrates that a similar staircase behavior emerges in the more complex maze (rectangular) task. In this case, the model’s behavior resembles a spreading “cloud” that progressively discovers contiguous areas of the maze. This is particularly noteworthy because the model is trained only on the final, fully solved maze configuration (shown in the second row of each column of Figure 23).
In both the cycles and maze tasks, we observe a consistent pattern of the model first solving easier, more local aspects of the problem before progressively tackling more global structures. This aligns with the theoretical understanding of the staircase effect in learning sparse Boolean functions, which is now demonstrated in the visual domain.